# On dense subalgebras of the singular ideal in groupoid C*-algebras

### v1

@@ 4. Ample Groupoids
Let $G$ be an étale groupoid, and let $f\in\mathscr{C}_{c}(G)$. Let $x\in G^{(0)}$, and assume that $f(g)=0$ for all $g\in G_{x}$. Then $f$ is continuous at each $g\in G_{x}$.
@@
Observations similar to the following lemma have appeared in both [4, Theorem 4.7] and [16, Proposition 5.19]. Our proof uses the same techniques as those developed in the aforementioned papers.
@@
Let $G$ be an étale groupoid. Fix $f\in J\cap\mathscr{C}_{c}(G)$ and $x_{0}\in G^{(0)}$. Let $g_{1},\dotsc,g_{n}\in G_{x_{0}}$ be all points in $G_{x_{0}}$ on which $f$ is non-zero, and let $U_{1},\dotsc,U_{n}$ be open bisections such that $U_{i}$ contains $g_{i}$ for each $1\leq i\leq n$. Set $V\coloneq\bigcap_{i=1}^{n}s(U_{i})$. Then there exists an open subset $W\subseteq V$ containing $x_{0}$ such that
$f^{\psi}\coloneq\sum_{i=1}^{n}f(g_{i})(\psi\circ s|_{U_{i}})$ (4.1)
satisfies $f^{\psi}\in J$ for any $\psi\in C_{c}(W)$.
@@
Without loss of generality, we assume $s(U_{i})=V$ for all $1\leq i\leq n$. For each open set $W\subseteq V$, let $S^{W}_{i}\coloneq U_{i}\cap s^{-1}(W)$ for $1\leq i\leq n$. For $I\subseteq\{1,\dotsc,n\}$, set $\check{S}^{W}_{I}\coloneq(\bigcap_{i\in I}S^{W}_{i})\setminus(\bigcup_{j\notin I}S^{W}_{j})$ as well as $W_{I}\coloneq\text{int}\Big(\overline{s(\check{S}^{W}_{I})}^{G^{(0)}}\Big)$, where $\overline{(\,\cdot\,)}^{G^{(0)}}$ denotes closure in $G^{(0)}$. Our first claim is that there exists an open set $W\subseteq V$ containing $x_{0}$ such that
$W_{I}\neq\emptyset\text{ implies that }x_{0}\in\overline{W_{I}}^{G^{(0)}}\text{ for all }I\subseteq\{1,\dotsc,n\}.$ (4.2)
If $W=V$ satisfies (4.2), then we are done. Otherwise, choose a regular open subset $W$ of $G^{(0)}$ (i.e., $W=\text{int}\big(\overline{W}^{G^{(0)}}\big)$) such that $x_{0}\in W\subseteq V$ and $W\cap\overline{V_{I}}^{G^{(0)}}=\emptyset$ for all $I\subseteq\{1,\dotsc,n\}$ with $V_{I}\neq\emptyset$ and $x_{0}\notin\overline{V_{I}}^{G^{(0)}}$. Note that $\check{S}^{W}_{I}=\check{S}^{V}_{I}\cap s^{-1}(W)$, and therefore $s(\check{S}^{W}_{I})=s(\check{S}^{V}_{I})\cap W$. Since $W$ is open, we have $W_{I}=V_{I}\cap W$, and therefore $\overline{W_{I}}^{G^{(0)}}\cap W=\overline{V_{I}}^{G^{(0)}}\cap W$. Also $W_{I}=\text{int}\Big(\overline{s(\check{S}^{W}_{I})}^{G^{(0)}}\Big)\subseteq\text{int}\Big(\overline{s(\check{S}^{V}_{I})}^{G^{(0)}}\cap\overline{W}^{G^{(0)}}\Big)\subseteq V_{I}\cap\text{int}\big(\overline{W}^{G^{(0)}}\big)=V_{I}\cap W=\emptyset$ whenever $x_{0}\notin\overline{V_{I}}^{G^{(0)}}$. This completes the proof of the claim.
Define $\mathcal{I}\coloneq\{I\subseteq\{1,\dotsc,n\}\colon W_{I}\neq\emptyset\}$. Our second claim is that
$\sum_{i\in I}f(g_{i})=0\text{ for all }I\in\mathcal{I}.$ (4.3)
By [4, Lemma 4.1(ii)], $\overline{s({\rm supp}^{\circ}(f))}^{G^{(0)}}$ has empty interior. Let $I\in\mathcal{I}$, and set $\check{W}_{I}\coloneq W_{I}\setminus\overline{s({\rm supp}^{\circ}(f))}^{G^{(0)}}$. Then, $\check{W}_{I}$ is open and dense in $W_{I}$. Therefore $x_{0}$ lies in the closure of $\check{W}_{I}$ in $G^{(0)}$ by (4.2). As $s(\check{S}_{I})$ is dense in $W_{I}$ and because $\check{W}_{I}$ is open, we conclude that there is a net $(x_{\beta})$ in $\check{W}_{I}\cap s(\check{S}_{I})$ converging to $x_{0}$ in $G^{(0)}$. Let $g_{\beta}$ be the unique element of $\check{S}_{I}$ such that $s(g_{\beta})=x_{\beta}$. By passing to a subnet if necessary, we may assume that $(g_{\beta})$ converges to some $\bm{g}$ in $\tilde{G}$ (i.e., in the Fell topology). By construction, we have that $\bm{g}\cap\{g_{1},\dotsc,g_{n}\}=\{g_{i}\colon i\in I\}$. Therefore,
$\lim_{\beta}f(g_{\beta})=\lim_{\beta}\mathfrak{i}(f)(\iota(g_{\beta}))=\mathfrak{i}(f)(\bm{g})=\sum_{i\in I}f(g_{i}),$
where the third equality used (1.2). Now $s(g_{\beta})\in\check{W}_{I}$ implies that $g_{\beta}\notin{\rm supp}^{\circ}(f)$, so that $f(g_{\beta})=0$. We conclude that $\sum_{i\in I}f(g_{i})=0$, proving (4.3).
Take arbitrary $\psi\in C_{c}(W)$, and let $f^{\psi}$ be as in (4.1). We show that $f^{\psi}\in J$. For $I\in\mathcal{I}$ and $g\in\check{S}^{W}_{I}$ we have $f^{\psi}(g)=\sum_{i\in I}f(g_{i})\psi(s(g))=0$ using (4.3). It follows that ${\rm supp}^{\circ}(f^{\psi})\subseteq\bigcup_{I\,\in\,\mathcal{P}(\{1,\dotsc,n\})\setminus\mathcal{I}}\check{S}^{W}_{I}$ and therefore $s({\rm supp}^{\circ}(f^{\psi}))\subseteq\bigcup_{I\,\in\,\mathcal{P}(\{1,\dotsc,n\})\setminus\mathcal{I}}s(\check{S}^{W}_{I})$, where $\mathcal{P}$ denotes power set. Hence $s({\rm supp}^{\circ}(f^{\psi}))$ has empty interior because it is contained in a finite union of nowhere dense sets. This shows that $f^{\psi}\in J$, as desired. ∎
@@
Let $f\in J$, $x_{0}\in G^{(0)}$ and assume that there are only finitely many points $g_{1},\dotsc,g_{n}$ in the isotropy group $G^{x_{0}}_{x_{0}}$ on which $f$ non-zero. Let $U_{1},\dotsc,U_{n},V$ be as in the statement of Lemma 4.2. If in addition we assume that the isotropy group $G^{x_{0}}_{x_{0}}$ is amenable, then the conclusion of Lemma 4.2 still holds. That is, there exists an open set $W\subseteq V$ containing $x_{0}$ such that $f^{\psi}\in J$ for any $\psi\in C_{c}(W)$, where $f^{\psi}$ is defined as in (4.1). If the isotropy group $G^{x_{0}}_{x_{0}}$ is non-amenable, then there may exist no such $W$.
We now prove the main result of this section. Recall that an ample groupoid is an étale groupoid with a basis of compact open bisections. If $G$ is an ample groupoid, the Steinberg algebra $\mathbb{C}G$ over the complex numbers is defined in [26] as
$\mathbb{C}G\coloneq\text{span}\left\{\mathbbm{1}_{U}\colon U\subseteq G\text{ a compact open bisection}\right\}$
The algebraic singular ideal is the intersection $J_{\mathbb{C}}\coloneq J\cap\mathbb{C}G$. Since $\mathbb{C}G\subseteq\mathscr{C}_{c}(G)$ we have $J_{\mathbb{C}}\subseteq J\cap\mathscr{C}_{c}(G)$. For $f\in\mathscr{C}_{c}(G)$, define
$\left\|f\right\|_{I}\coloneq\max\left\{\sup_{x\in G^{(0)}}\sum_{g\in G_{x}}\left|f(g)\right|,\sup_{x\in G^{(0)}}\sum_{g\in G^{x}}\left|f(g)\right|\right\}.$

### v2

@@ 5.1. Preliminary lemmas
Let $G$ be an étale groupoid. Fix $f\in\mathscr{C}_{c}(G)$, $x\in G^{(0)}$, and assume that $f(g)=0$ for all $g\in G_{x}$. Then $f$ is continuous at each $g\in G_{x}$.
@@
The following results isolate observations from [19, Proposition 5.19]. We follow the same proofs (see also [5, Theorem 4.7]). Given $x\in G^{(0)}$, recall that a subgroup $X\subseteq G^{x}_{x}$ belongs to $\mathcal{X}(x)$ if and only if there exists a net $(x_{\alpha})\subset C$ of (Hausdorff) units whose set of limit points is precisely $X$, and such that every subnet of $(x_{\alpha})$ has its limit points contained in $X$ (see Definition 1.3).
@@
Let $G$ be an étale groupoid that is covered by countably many open bisections. Fix $x\in G^{(0)}$, and let $g_{1},\dots,g_{n}\in G_{x}$ be distinct points.
There exist open bisections $U_{1},\dots,U_{n}\subseteq G$ satisfying the following conditions:
  1. (i)
    $g_{i}\in U_{i}$ for all $i\in\{1,\dots,n\}$;
  2. (ii)
    $s(U_{i})=s(U_{j})$ for all $i,j\in\{1,\dots,n\}$;
  3. (iii)
    $W_{I}=\emptyset$ whenever $I\subseteq\{1,\dotsc,n\}$ is not of the form
    (5.1) $\{g_{i}\colon i\in I\}=kX\cap\{g_{1},\dots,g_{n}\}$
    for some $k\in G_{x}$ and $X\in\mathcal{X}(x)$.
Here $W_{I}\coloneq s\big(\bigcap_{i\in I}U_{i}\setminus\bigcup_{j\notin I}U_{j}\big)\cap C$.
Note that for any small open set $U\subseteq G^{(0)}$ containing $x$, the cutdowns $U_{i}^{\prime}\coloneq U_{i}\cap s^{-1}(U)$ also satisfy (i) – (iii).
@@
It is clear that there exist open bisections $V_{1},\dots,V_{n}\subseteq G$ satisfying the first two conditions. Write $V\coloneq s(V_{1})$ and $V_{I}\coloneq s\big(\bigcap_{i\in I}U_{i}\setminus\bigcup_{j\notin I}U_{j}\big)\cap C$ for subsets $I\subseteq\{1,\dots,n\}$. Choose an open subset $W\subseteq G^{(0)}$ such that $x\in W\subseteq V$ and $W\cap\overline{V_{I}}^{G^{(0)}}=\emptyset$ for all $I\subseteq\{1,\dotsc,n\}$ with $x\notin\overline{V_{I}}^{G^{(0)}}$. Define $U_{i}\coloneq V_{i}\cap s^{-1}(W)$ for $i\in\{1,\dots,n\}$. It is clear that the open bisections $U_{1},\dots,U_{n}\subseteq G$ satisfy conditions (i) and (ii). We claim that they also satisfy condition (iii).
Let $I\subseteq\{1,\dots,n\}$ be such that $W_{I}\neq\emptyset$. We claim that $I$ is of the form in (5.1). It is clear that $W_{I}=W\cap V_{I}$ and hence $x\in\overline{W_{I}}^{G^{(0)}}$ by the construction of $W$. Let $(x_{\alpha})$ be a net in $W_{I}$ converging to $x$. By passing to a subnet if necessary we can assume that the net $\iota(x_{\alpha})$ converges in the Hausdorff cover groupoid $\tilde{G}$. Let $X\in\tilde{G}^{(0)}_{\rm{ess}}$ denote the limit. Note that $X\in\mathcal{X}(x)$ by definition. Also, by definition of the Fell topology, the set of limit points of the net $(x_{\alpha})$ is precisely $X$. Now, let $g_{\alpha}\in\bigcap_{i\in I}U_{i}\setminus\bigcup_{j\notin I}U_{j}$ be the unique elements satisfying $s(g_{\alpha})=x_{\alpha}$. Continuity of multiplication on the groupoid implies that the net $(g_{\alpha})$ has set of limit points $g_{i_{0}}X$, where $i_{0}$ is any element of $I$. The unit $x$ belongs to $X$, and hence $\{g_{i}\colon i\in I\}\subseteq g_{i_{0}}X$. On the other hand, if $j\notin I$ then $g_{\alpha}\notin U_{j}$ and hence the net $(g_{\alpha})$ does not converge to $g_{j}$. Therefore $g_{j}\notin g_{i_{0}}X$, and it follows that $\{g_{i}\colon i\in I\}=g_{i_{0}}X\cap\{g_{1},\dots,g_{n}\}$ for any $i_{0}\in I$. This completes the proof. ∎
The next proposition is a slight generalisation of the construction of singular elements found in [19, Proposition 5.19]. The proof is exactly the same, but we record it for completeness.
Proposition 5.3.
Let $G$ be an étale groupoid that is covered by countably many open bisections. Fix $x\in G^{(0)}$, and suppose $b\in\mathbb{C}[G^{x}_{x}]$ belongs to $\ker(\lambda_{G^{x}_{x}/X})$ for all $X\in\mathcal{X}(x)$. Let $\{g_{1},\dots,g_{n}\}\subseteq G^{x}_{x}$ denote the support of $b$, let $h\in G_{x}$, and let $U_{1},\dots,U_{n}$ be open bisections containing $hg_{1},\dots,hg_{n}$ and satisfying the conditions of Lemma 5.2. Then
(5.2) $f^{\psi}\coloneq\sum_{i=1}^{n}b(g_{i})(\psi\circ s|_{U_{i}})$
belongs to $J\cap\mathscr{C}_{c}(G)$ for any $\psi\in C_{c}\big(s(U_{i})\big)$ (note that this does not depend on $i$).
Proof.
For $I\subseteq\{1,\dots,n\}$ set $\check{S}_{I}=(\bigcap_{i\in I}U_{i})\setminus(\bigcup_{j\notin I}U_{j})$ and $W_{I}=s(\check{S}_{I})\cap C$, as in Lemma 5.2. Take arbitrary $\psi\in C_{c}\big(s(U_{i})\big)$, and let $f^{\psi}$ be as in (5.2). If $W_{I}\neq\emptyset$, Lemma 5.2(iii) ensures that $\{hg_{i}\colon i\in I\}=kX\cap\{hg_{1},\dots,hg_{n}\}$ for some $k\in G_{x}$ and $X\in\mathcal{X}(x)$. Clearly $h^{-1}k\in G^{x}_{x}$. Therefore, for $g\in\check{S}_{I}$ we have
$f^{\psi}(g)=\sum_{i\in I}b(g_{i})\psi(s(g))=\sum_{t\in h^{-1}kX}b(t)\psi(s(g))=0$
since $b\in\ker(\lambda_{G^{x}_{x}/X})$. It follows that any $g\in{\rm supp}^{\circ}(f^{\psi})$ belongs to $\check{S}_{I}$ for some $I\subseteq\{1,\dots,n\}$ with $W_{I}=\emptyset$, and hence $s\big({\rm supp}^{\circ}(f^{\psi})\big)\subseteq G^{(0)}\setminus C$. This implies that $f^{\psi}\in J$ by [22, Proposition 7.18], as desired. ∎
$\bullet$
5.2. Building blocks for $J\cap\mathscr{C}_{c}(G)$
Let $G$ be an étale groupoid that is covered by countably many open bisections. Each building block is assembled from the data in a tuple $\Xi\coloneq(x,h,b,U_{1},\dots,U_{n},\psi)$ where
  • $x\in G^{(0)}$;
  • $h\in G_{x}$;
  • $b\in\mathbb{C}[G^{x}_{x}]$ belongs to $\ker(\lambda_{G^{x}_{x}/X})$ for all $X\in\mathcal{X}(x)$, and has support $\{g_{1},\dots,g_{n}\}$;
  • $U_{1},\dots,U_{n}\subseteq G$ are open bisections satisfying the conditions of Lemma 5.2. That is, they have equal source, contain the elements $hg_{1},\dots,hg_{n}\in G_{x}$, and are such that $W_{I}=\emptyset$ whenever $I\subseteq\{1,\dotsc,n\}$ is not of the form
    $\{hg_{i}\colon i\in I\}=kX\cap\{hg_{1},\dots,hg_{n}\}\quad\quad\text{for some }k\in G_{x},X\in\mathcal{X}(x).$
    Here $W_{I}\coloneq s\big(\bigcap_{i\in I}U_{i}\setminus\bigcup_{j\notin I}U_{j}\big)\cap C$.
  • $\psi\in C_{c}\big(s(U_{i})\big)$ (note that this does not depend on $i$).
Given such a tuple $\Xi$ we define
(5.3) $f^{\Xi}\coloneq\sum_{i=1}^{n}b(g_{i})(\psi\circ s|_{U_{i}}).$
By Proposition 5.3 we have $f^{\Xi}\in J\cap\mathscr{C}_{c}(G)$. Let $\mathcal{E}\subseteq J\cap\mathscr{C}_{c}(G)$ denote the linear span of elements of the form $f^{\Xi}$ from (5.3), where we let $\Xi$ range over all valid tuples. We will see that $\mathcal{E}$ is dense in $J\cap\mathscr{C}_{c}(G)$.
Lemma 5.4.
Let $G$ be an étale groupoid that is covered by countably many open bisections. Fix $f\in J\cap\mathscr{C}_{c}(G)$, $x\in G^{(0)}$, and let $g_{1},\dotsc,g_{n}\in G_{x}$ be all elements of $G_{x}$ on which $f$ is non-zero. There exist open bisections $U_{1},\dots,U_{n}$ with equal source, containing $g_{1},\dots,g_{n}$ and such that
$f^{\psi}\coloneq\sum_{i=1}^{n}f(g_{i})(\psi\circ s|_{U_{i}})$
belongs to $\mathcal{E}$ for any $\psi\in C_{c}\big(s(U_{i})\big)$ (note that this does not depend on $i$).
Proof.
Let $U_{1},\dots,U_{n}$ be open bisections satisfying the conditions of Lemma 5.2 and such that $U_{i_{1}}$ and $U_{i_{2}}$ are disjoint whenever $r(g_{i_{1}})\neq r(g_{i_{2}})$. Partition the set $\{g_{1},\dots,g_{n}\}$ according to the images under the range map $r$ and write $\{g_{1},\dots,g_{n}\}=\bigsqcup_{j=1}^{m}\{g^{j}_{1},\dots,g^{j}_{k_{j}}\}$, $\{U_{1},\dots,U_{n}\}=\bigsqcup_{j=1}^{m}\{U^{j}_{1},\dots,U^{j}_{k_{j}}\}$, where $g_{i}^{j}\in U_{i}^{j}$. Fix $\psi\in C_{c}\big(s(U_{i})\big)$. For each $j\in\{1,\dots,m\}$ define
$f^{j}\coloneq\sum_{i=1}^{k_{j}}f\big(g^{j}_{i}\big)\big(\psi\circ s|_{U^{j}_{i}}\big)$
so that $f^{\psi}=\sum_{j=1}^{m}f^{j}$. For each $j$, we show that $f^{j}=f^{\Xi_{j}}$ for some valid tuple $\Xi_{j}$, where $f^{\Xi_{j}}$ is as in (5.3). The lemma then follows.
Let $V_{j}$ be an open bisection containing ${(g^{j}_{1})}^{-1}$, and let $\varphi\in C_{c}(V_{j})$ be such that $\varphi\big(({g^{j}_{1}}\big)^{-1})=1$. Define $b_{j}\coloneq\eta_{x}(\varphi*f)$, where $\eta_{x}\colon\mathscr{C}_{c}(G)\to\mathbb{C}[G^{x}_{x}]$ is the restriction map. Since $\varphi*f\in J\cap\mathscr{C}_{c}(G)$, it follows from [19, Proposition 5.19] that $b_{j}$ belongs to $\ker(\lambda_{G^{x}_{x}/X})$ for all $X\in\mathcal{X}(x)$. Then, since $f(g_{i}^{j})=b_{j}\big({(g^{j}_{1})}^{-1}g^{j}_{i}\big)$ for all $i=1,\dots,k_{j}$, we have $f^{j}=f^{\Xi_{j}}$ where $\Xi_{j}\coloneq\big(x,g_{1}^{j},b_{j},U^{j}_{1},\dots,U^{j}_{k_{j}},\psi\big)$. The initial assumptions on the bisections $U_{1},\dots,U_{n}$ ensure that $\Xi_{j}$ is a valid tuple. ∎
For $f\in\mathscr{C}_{c}(G)$, define
(5.4) $\left\|f\right\|_{I}\coloneq\max\left\{\sup_{x\in G^{(0)}}\sum_{g\in G_{x}}\left|f(g)\right|,\sup_{x\in G^{(0)}}\sum_{g\in G^{x}}\left|f(g)\right|\right\},$
following [16].
Theorem 5.5.
Let $G$ be an étale groupoid that is covered by countably many open bisections. Then $\mathcal{E}$ is dense in $J\cap\mathscr{C}_{c}(G)$ with respect to the norm $\left\|\cdot\right\|_{I}$.
The norm $\left\|\cdot\right\|_{I}$ dominates the full norm $\left\|\cdot\right\|_{C^{*}(G)}$ and hence any $C^{*}$-norm. Therefore, $\mathcal{E}$ is dense in $J\cap\mathscr{C}_{c}(G)$ with respect to any $C^{*}$-norm.
Proof.
Let $f\in J\cap\mathscr{C}_{c}(G)$. Take a compact set $K\subseteq G$ with ${\rm supp}^{\circ}(f)$ contained in its interior. Whenever $\dot{f}\in\mathscr{C}_{c}(G)$ satisfies ${\rm supp}^{\circ}(\dot{f})\subseteq K$ we have $\left\|\dot{f}\right\|_{I}\leq C_{K}\left\|\dot{f}\right\|_{\infty}$ where $\left\|\cdot\right\|_{\infty}$ denotes the supremum norm, and $C_{K}$ is a constant depending only on $K$. Therefore, it suffices to find, for each $\varepsilon>0$, a function $f^{\prime}\in\mathcal{E}$ satisfying ${\rm supp}^{\circ}(f^{\prime})\subseteq K$ and $\left\|f-f^{\prime}\right\|_{\infty}\leq\varepsilon$.
Fix $\varepsilon>0$. Take $x\in s(K)$, and let $g_{1},\dotsc,g_{n}\in G_{x}$ be all elements of $G_{x}$ on which $f$ is non-zero. Let $U_{1},\dots,U_{n}\subseteq K$ be open bisections as in Lemma 5.4, and define $U_{x}\coloneq s(U_{i})$ (note that this does not depend on $i$). In particular $x\in U_{x}$. Then
$f^{\psi}\coloneq\sum_{i=1}^{n}f(g_{i})(\psi\circ s|_{U_{i}})$
belongs to $\mathcal{E}$ for any $\psi\in C_{c}(U_{x})$. Fix some $\psi_{x}\in C_{c}(U_{x})$ with $\psi_{x}(x)=1$. Clearly ${\rm supp}^{\circ}(f^{\psi_{x}})\subseteq K$. We have $f|_{G_{x}}=f^{\psi_{x}}|_{G_{x}}$ and ${\rm supp}^{\circ}(f-f^{\psi_{x}})\subseteq K$, so by Lemma 5.1 there exists an open neighbourhood $W_{x}\subseteq U_{x}$ of $x$ such that $\left|f(g)-f^{\psi_{x}}(g)\right|\leq\varepsilon$ whenever $s(g)\in W_{x}$. The sets $\{W_{x}\}_{x\in s(K)}$ form an open cover for the compact set $s(K)$, so select a finite subcover $W_{x_{1}},\dotsc,W_{x_{m}}$. Let $\varphi_{1},\dots,\varphi_{m}\in C_{c}\big(G^{(0)}\big)$ be a partition of unity for $s(K)$ subordinate to this cover, and define
$f^{\prime}\coloneq\sum_{j=1}^{m}(\varphi_{j}\circ s)f^{\psi_{x_{j}}}=\sum_{j=1}^{m}f^{\varphi_{j}.\psi_{x_{j}}}.$
Then $f^{\prime}\in\mathcal{E}$ satisfies ${\rm supp}^{\circ}(f^{\prime})\subseteq K$ and $\left\|f-f^{\prime}\right\|_{\infty}\leq\varepsilon$, as required. ∎
@@ 5.2. Building blocks for $J\cap\mathscr{C}_{c}(G)$
The construction in this subsection generalises to groupoids $G$ that are not covered by countably many open bisections. For an open subgroupoid $H\subseteq G$ that is covered by countably many open bisections, define $\mathcal{E}_{H}$ to be the linear span of functions of the form $f^{\Xi}$ in (5.3) for tuples $\Xi$ in $H$. Then, define $\mathcal{E}\subseteq J\cap\mathscr{C}_{c}(G)$ to be the subspace generated by all the $\mathcal{E}_{H}$. With this modification, the conclusion of Theorem 5.5 holds for any étale groupoid.
6. Ample Groupoids
In this section we describe an explicit spanning set for the algebraic singular ideal $J_{\mathbb{C}}$ of an ample groupoid. The construction is similar to that in Subsection 5.2. In Subsection 6.2 we prove that the algebraic singular ideal $J_{\mathbb{C}}$ is dense in $J\cap\mathscr{C}_{c}(G)$ for any ample groupoid.
Recall that an ample groupoid is an étale groupoid with a basis of compact open bisections. If $G$ is an ample groupoid, the Steinberg algebra $\mathbb{C}G$ over the complex numbers is defined in [33] as
$\mathbb{C}G\coloneq\text{span}\left\{\mathbbm{1}_{U}\colon U\subseteq G\text{ a compact open bisection}\right\}.$
The algebraic singular ideal is the intersection $J_{\mathbb{C}}\coloneq J\cap\mathbb{C}G$. Since $\mathbb{C}G\subseteq\mathscr{C}_{c}(G)$, we have $J_{\mathbb{C}}\subseteq J\cap\mathscr{C}_{c}(G)$.
$\bullet$
6.1. Building blocks for $J_{\mathbb{C}}$
Let $G$ be an ample groupoid that is covered by countably many open bisections. Each building block is assembled from the data in a tuple $\Theta\coloneq(x,h,b,U_{1},\dots,U_{n})$ where
  • $x\in G^{(0)}$;
  • $h\in G_{x}$;
  • $b\in\mathbb{C}[G^{x}_{x}]$ belongs to $\ker(\lambda_{G^{x}_{x}/X})$ for all $X\in\mathcal{X}(x)$, and has support $\{g_{1},\dots,g_{n}\}$;
  • $U_{1},\dots,U_{n}\subseteq G$ are compact open bisections satisfying the conditions of Lemma 5.2. That is, they have equal source, contain the elements $hg_{1},\dots,hg_{n}\in G_{x}$, and are such that $W_{I}=\emptyset$ whenever $I\subseteq\{1,\dotsc,n\}$ is not of the form
    $\{hg_{i}\colon i\in I\}=kX\cap\{hg_{1},\dots,hg_{n}\}\quad\quad\text{for some }k\in G_{x},X\in\mathcal{X}(x).$
    Here $W_{I}\coloneq s\big(\bigcap_{i\in I}U_{i}\setminus\bigcup_{j\notin I}U_{j}\big)\cap C$.
Given such a tuple $\Theta$ we define
(6.1) $f^{\Theta}\coloneq\sum_{i=1}^{n}b(g_{i})\mathbbm{1}_{U_{i}}.$
Then $f^{\Theta}\in J_{\mathbb{C}}$ by Proposition 5.3.